Find a linear differential operator that annihilates the given function. (Use D for the differential operator.)

7x − sin(x) + 20 cos(5x)


1
Expert's answer
2021-03-30T10:56:19-0400

The differential operator "(D-\alpha)^n" annihilates the functions:


"e^{\alpha x},xe^{\alpha x},x^2e^{\alpha x}, ..., x^{n-1}e^{\alpha x}"

Then the differential operator "D^2" annihilates the function "x."


We know that the differential operator "D^2-2\alpha D+(\alpha ^2+\beta ^2)" annihilates "e^{\alpha x}\cos(\beta x)" or "e^{\alpha x}\sin(\beta x)."

Thus the differential operator "D^2+1" annihilates the function "\sin(x),"

the differential operator "D^2+25" annihilates the function "\cos(5x)."


A linear differential operator that annihilates the given function is


"D^2(D^2+1)(D^2+25)"